Let X1. . . . Xn be i.i.d Uniform over the interval (θ, θ + 1].Show that X(1)+X(n) )/2- 1/2 is also an unbiased estimator of θ, whereX(1) is the minimum order statistic and X(n) is the maximum order statistic. If X - 1/2 is also an unbiased estimator of θ which of the two estimators would you prefer to use.
Answer #1

Answer:-

Given That:-

Let X1. . . . Xn be i.i.d Uniform over the interval (θ, θ + 1].Show that X(1)+X(n) )/2- 1/2 is also an unbiased estimator of θ, whereX(1) is the minimum order statistic and X(n) is the maximum order statistic. If X - 1/2 is also an unbiased estimator of θ which of the two estimators would you prefer to use.

X_{1}, X_{2}, ...X_{n}\sim U(\theta, \theta+1)

Let

X_{i} \sim U(\theta, \theta+1),   

\theta < X_{i} < \theta + 1

\Rightarrow \theta < X_{i} - \theta < 1

X_{i} - \theta \sim U(0, \theta+1)\Rightarrow 0 < X_{i} - \theta < 1

Put

Y_{i} = X_{i} - \theta

then

Y_{i} \sim U(0, 1),0 < y_{i} < 1

Result:-

If  y_{(1)}, y_{(2)}, .... y_{(n)} are order statics of  y_{(1)}, y_{(2)}, .... y_{(n)}, y_{i} \sim U(0, 1)

then r th order statistic  y_{(r)} \sim beta\, \, \, (r, n-r+1)

Therefore  y_{(r)} \sim beta\, \, \, (r, n-r+1)

Therefore y_{(1)}\sim \beta (1, n)

y_{(n)}\sim \beta (n, 1)

If X \sim beta\, \, \, (\alpha, \beta)

E(X) = \frac{\alpha}{\alpha + \beta}

Var(X) = \frac{\alpha \beta}{(\alpha + \beta)^{2}(\alpha + \beta + 1)}

\therefore E(y_{(1)}) = \frac{1}{n + 1}

Put

y_{(1)} = X_{(1)} - \theta

\Rightarrow E(X_{(1)} - \theta) = \frac{1}{n + 1}

\Rightarrow E(X_{(1)} = \frac{1}{n + 1} + \theta

E(y_{(n)}) = \frac{n}{n+1} \Rightarrow E(X_{(n)} - \theta) = \frac{n}{n+1}

\Rightarrow E(X_{(n)}) = \frac{n}{n+1} + \theta

\Rightarrow \frac{X_{(1)} + X_{(n)}}{2} - \frac{1}{2}

\Rightarrow E\left ( \frac{X_{(1)} + X_{(n)}}{2} - \frac{1}{2} \right ) = \frac{1}{2} \left [ E(X_{(1)} + E(X_{(n)})) \right ] - \frac{1}{2}

= \frac{1}{2}\left [ \frac{1}{n+1} + \theta + \frac{n}{n+1}+\theta \right ] - \left [ \frac{1}{2} \right ]

\Rightarrow \frac{1}{2}\left [ \frac{1}{n+1} [n+1] + 2\theta \right ] - \left [ \frac{1}{2} \right ]

\Rightarrow \frac{1}{2} + \theta - \frac{1}{2} = \theta

\therefore \frac{X_{(1)} + X_{(n)}}{2} - \frac{1}{2} is an unbaised estimation for \theta

X\sim U(\theta, \theta + 1)

E(X) = \frac{\theta + \theta + 1}{2}

= \frac{2\theta + 1}{2}

E(X - \frac{1}{2})= \frac{2\theta + 1}{2} = \theta

From both of these Statistic will choose the statistic which have minimum variance.

\therefore Var\left ( \frac{X_{(1)} + X_{(n)}}{2} - \frac{1}{2} \right ) = \frac{1}{4}\left [ Var(X_{(1)}) + Var(X_{(n)}) \right ]

  = \frac{1}{4}\left [ \frac{n}{(n + 1)^{2}(n + 2)} + \frac{n}{(n + 1)^{2}(n + 2)} \right ]

= \frac{1}{2}\left ( \frac{n}{(n+1)^{2}(n + 2)} \right )

\therefore Var(X - \frac{1}{2}) = \frac{1}{12}

\therefore Var\left ( \frac{X_{(1)} + X_{(n)}}{2} - \frac{1}{2} \right ) have minimum variance will prefer these estimator.

Plz like it...,